OpenStudy (anonymous):

Hi! Can someone help me figure out how to solve this : The height in feet of an object is given by y = -16t2 + 80t + 96, where y is the height and t is the time in seconds that the object is in the air. After how much time does the object hit the ground?

OpenStudy (solomonzelman):

y is the position of the subject, and t is the number of seconds. You need find the time when it hits the ground, i.e. to find the value(s) of t when y=0.

OpenStudy (solomonzelman):

Set \(\large\color{slate}{ 0=-16t^2 + 80t + 96 }\) and solve for t.

OpenStudy (anonymous):

should I use the distance formula? meh I'm confused lol

OpenStudy (anonymous):

not when you are solving a quadratic equation, no

OpenStudy (anonymous):

forget about the words, just set it equal to zero and solve what @SolomonZelman wrote

OpenStudy (anonymous):

oh ok I get it thank you! 'll post the answer after I solve it.

OpenStudy (misty1212):

HI!!

OpenStudy (solomonzelman):

sure:)

OpenStudy (anonymous):

Hi ( :

OpenStudy (misty1212):

divide all by \(-16\) first, it will make it loads easier

OpenStudy (anonymous):

ah I can't see what you typed I have a "math processing error"

OpenStudy (solomonzelman):

she said that dividing the quadratic equation by -16 would make it a lot easier to solve.(divide every term in the equation by -16)

OpenStudy (anonymous):

refresh your browser, usually works

OpenStudy (solomonzelman):

btw, math process error is something that I often have. Switching browsers also works (in fact, more than refreshing, for me).

OpenStudy (anonymous):

does -2 or -3 sound right?

OpenStudy (anonymous):

as the solution set

OpenStudy (misty1212):

no

OpenStudy (misty1212):

\[0 = -16 (t-6) (t+1)\]

OpenStudy (solomonzelman):

no

OpenStudy (anonymous):

thanks @misty1212 I'm working on it

OpenStudy (anonymous):

1/2 or -1/8?

OpenStudy (anonymous):

@SolomonZelman when I divided by -16 I got -t^2 + -5t + -6 = 0

OpenStudy (anonymous):

solution set of 6,-1

OpenStudy (anonymous):

but when I tried to plug it in it didn't work out

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle 0=-16t^2 + 80t + 96}\) \(\large\color{slate}{\displaystyle 0=-16(t^2 -5t -6)}\) \(\large\color{slate}{\displaystyle 0=-16(t+1)(x-6)}\)

OpenStudy (solomonzelman):

for t, in a quadratic equation you DO get -1 and 6, but remember, that you are finding "the seconds' because t is seconds.

OpenStudy (solomonzelman):

And you can't have -1 seconds.

OpenStudy (solomonzelman):

so, you only include the positive solution for t.

OpenStudy (anonymous):

True, thank you! :D

OpenStudy (solomonzelman):

yw